# Enhancement of axial anomaly effects in hot two-color QCD: FRG approach in the linear sigma model

### v1

@@ IV.3 Enhancement of anomaly effects at finite $T$
For any values of $\mathring{M}_{\eta}$, all the coefficients exhibit significant increase at finite temperature, that is, the anomaly enhancement in hot medium is predicted to be similar in QC2D to the case of (three-color) QCD Fejos and Hosaka (2016).
Figure 4: $T$ dependencies of the mean field $\sigma_{0}$ with $\mathring{M}_{\eta}=0.3$, $0.5$ and $0.95$ GeV.
Figure 5: $T$ dependencies of the normalized topological susceptibilities $\chi_{\rm top}$ (solid) and $\tilde{\chi}_{\rm top}$ (dashed) defined in Eqs. (57) and (58), respectively, with $\mathring{M}_{\eta}=0.3$, $0.5$ and $0.95$ GeV.
The enhancement of the quadratic anomaly coefficient $a$ is milder than the quartic ones, $c_{1}$ and $c_{2}$, where a larger $\mathring{M}_{\eta}$ leads to a smaller increase. For $c_{1}$ and $c_{2}$, the same ordering of the enhancement is seen up to $T\sim 0.2$ GeV, while at higher temperatures it is flipped. In general, the coefficients show a significant increase at low $T$, while at high $T$ the enhancements become tempered and begin to decrease. Note that the inversions occur when the chiral condensate starts to reduce (see Fig. 4). In addition, such non-monotonic behaviors appear at lower $T$ for smaller values of $\mathring{M}_{\eta}$. We also note that $a<0$, $c_{1}<0$ and $c_{2}>0$ at any $T$.
@@ IV.4 Topological susceptibility
In terms of the fermionic language, this quantity can be re-expressed through two-point functions of pseudoscalar operators $\bar{\psi}i\gamma_{5}\psi$ and $\bar{\psi}i\gamma_{5}\tau_{f}^{a}\psi$, by the virtue of axial Ward-Takahashi identities Gómez Nicola and Ruiz de Elvira (2016); Kawaguchi et al. (2021). Then, within the LSM, imposing matching conditions appropriately the topological susceptibility is evaluated as (for the detailed derivation, see Ref. Kawaguchi and Suenaga (2023))
@@
are pion and $\eta$-meson propagators in momentum space.
In the present work, $D_{\pi}$ and $D_{\eta}$ are always simply expressed by $D_{\pi}=i/(p^{2}-M_{\pi}^{2})$ and $D_{\eta}=i/(p^{2}-M_{\eta}^{2})$, respectively. Note that, their functional forms do not differ even at finite $T$, owing to the absence of diquark condensates Kawaguchi and Suenaga (2023). Hence, the susceptibility reads
@@
where we have defined the dimensionless topological susceptibility, $\chi_{\rm top}$, normalized by the constant $\mathring{M}_{\pi}^{2}\mathring{\sigma}_{0}^{2}$. The mass difference between the $\eta$ meson and the pion is generated solely by anomaly effects [see Eq. (80)], therefore, it is beneficial to define
@@
Depicted in Fig. 5 are the resulting $T$ dependencies of the normalized topological susceptibilities $\chi_{\rm top}$ and $\tilde{\chi}_{\rm top}$ with $\mathring{M}_{\eta}=0.3$, $0.5$ and $0.95$ GeV. Differences between the solid and dashed curves are always tiny, meaning that the anomaly enhancement effects do not generate sizable contributions to the topological susceptibility. Meanwhile, the dependence of the $\eta$ meson mass on the topological susceptibility is only manifest for larger $\mathring{M}_{\eta}$, which increases the magnitude of $\chi_{\rm top}$. Up to $T\sim 0.1$ GeV the susceptibilities are almost constant for any values of $\mathring{M}_{\eta}$, while above this temperature they start to reduce. This reduction is understood by the smooth chiral symmetry restoration exhibited in Fig. 4, since the $T$ dependence of the susceptibility (57) is governed by
@@
Figure 6: $k$ dependencies of the anomaly coefficients $c_{1,k}$ and $c_{2,k}$ evaluated by the full (solid) and truncated (dashed) analyses at $T=0$, with $\mathring{M}_{\eta}=0.5$ GeV. Note that the UV scale is set to be $\Lambda_{\rm UV}=1$ GeV.
@@ V Importance of baryonic contributions
In the previous section, modifications of the hadron masses and anomaly coefficients were examined. In this section, we shed light on the role of baryonic contributions in the FRG framework to gain insight into fluctuations specific to QC2D, where diquark baryons are also counted as hadrons. Here we concentrate on the case where $\mathring{M}_{\eta}=0.5$ GeV.
In order to study baryonic contributions, on top of the full solution, we also solve the flow equations containing solely mesonic contributions (e.g., $\partial_{k}m^{2}_{k}=\tilde{\partial}_{k}\hat{V}^{I_{1}}_{M,k}$ for $m^{2}_{k}$) with inputs of $\mathring{M}_{\pi}=0.14$ GeV, $\mathring{M}_{\eta}=0.5$ GeV, $\lambda_{1,\Lambda}=\lambda_{2,\Lambda}/4$ and $c_{1,\Lambda}=c_{2,\Lambda}=0$. The value of $\lambda_{1,\Lambda}$ is adjusted such that $\mathring{M}_{\sigma}\approx 0.18$ GeV. Then we compare the cases with and without baryons.
The parameters $m^{2}_{k}$, $a_{k}$ and $\lambda_{1,k}$ at $k=0$ are roughly determined by the masses of the pion, $\eta$ and $\sigma$ mesons, respectively, and hence those values in the truncated and in the full analysis do not differ much. On the other hand, $c_{1,k}$ and $c_{k,2}$ are simply fixed to vanish at $k=\Lambda_{\rm UV}$, and thus their scale evolutions are expected to be a useful measure to see the differences between the two approaches. Depicted in Fig. 6 are the resultant $k$ dependencies for $c_{1,k}$ and $c_{2,k}$ at $T=0$, evaluated separately for the full solution and that of containing only mesonic contributions. This figure clearly shows that the evolutions of $c_{1,k}$ and $c_{2,k}$ start to deviate in the two cases, soon after departing from $k=\Lambda_{\rm UV}$. As a result, at $k=0$, the coefficients containing both mesonic and baryonic contributions are significantly magnified compared to the truncated case, indicating the important role of baryonic fluctuations. Such enhancements can be simply understood by an increase in the degrees of freedom in the flows.
Finally, we also show in Fig. 7 the temperature dependencies of the anomaly coefficients at $k=0$ evaluated via the full and the truncated analyses. This figure also shows the importance of baryonic contributions at any temperature with regards to the anomaly coefficients.
@@ VI Conclusions
In this paper we have investigated modifications of $U(1)_{A}$ anomaly effects at finite temperature in QC2D with two flavors, by means of the FRG technique. We have adopted the LSM developed in Ref. Suenaga et al. (2023) to properly take into account fluctuations of the low-lying scalar and pseudoscalar mesons, as well as baryons, based on a linear representation of the Pauli-Gürsey $SU(4)$ symmetry.
As a result, we have found that, quartic couplings of the LSM need to be suppressed, in order to properly reproduce a reasonable reduction of the chiral condensate at finite temperature. Accordingly, mass differences between the chiral partners, $\sigma$ and $\pi$, or, $a_{0}$ and $\eta$, are always found to be tiny, regardless of the $U(1)$ axial anomaly strength, i.e., the mass of $\eta$ meson. At higher temperatures, mass degeneracies of the chiral partners in accordance with chiral symmetry restoration are realized. We note that masses of the $0^{+}$ and $0^{-}$ (anti)diquark baryons degenerate with those of $\pi$ and $a_{0}$, respectively, owing to the Pauli-Gürsey $SU(4)$ symmetry. The topological susceptibility at high temperature is found to be suppressed due to the reduction of chiral condensate.
Once we enter the finite-temperature regime, all of the couplings responsible for the $U(1)$ axial anomaly become magnified, and hence we have concluded that due to hadronic fluctuations, the axial anomaly is enhanced in a hot medium for QC2D. Similar anomaly enhancements were also discussed in three-color QCD analyses Fejos and Hosaka (2016). At high temperatures, the increase in the quartic couplings is gradually suppressed and eventually diminishes. Our results also confirm that baryonic fluctuations play a substantial role in the enhancements.
@@
We note that, while the $U(1)_{A}$ anomaly effects are enhanced at finite temperature within the hadronic model, LSM, it is known that at sufficiently high temperature and/or density, instantons melt due to the Debye screening of electric gluons, for which the $U(1)_{A}$ symmetry is effectively restored Gross et al. (1981); Rapp et al. (2000). In order to smoothly connect those two behaviors, it would be necessary to implement a crossover transition to quark matter, but it is beyond the scope of the present study.
@@
Our analysis can also be applied to the case of a finite chemical potential. In the low-temperature and high-density regime, the emergence of a baryon superfluid phase, characterized by the condensation of positive-parity diquarks, has been demonstrated in lattice QCD simulations Boz et al. (2020); Buividovich et al. (2020); Astrakhantsev et al. (2020); Iida et al. (2024); Braguta (2023) and model analyses Kogut et al. (2000); Ratti and Weise (2004); Sun et al. (2007); Suenaga et al. (2023); Suenaga (2025). In this phase, in particular, the possibility of enhanced $U(1)_{A}$ anomaly effects has been discussed by comparing the lattice data with LSM analyses of the hadron mass spectrum and topological susceptibilities  Suenaga et al. (2023); Kawaguchi and Suenaga (2023). Hence, investigating the fate of the $U(1)_{A}$ anomaly effects in the superfluid phase is considered an important direction for future work.
Finally, we provide some comments regarding diquarks. In three-color QCD, the chiral dynamics of diquarks is reflected in the spectrum of singly heavy baryons (SHBs), which consist of a heavy quark $Q$ and a light diquark, either [$qq$] or [$sq$] ($q=u,d$). In this context, it is suggested that the $U(1)_{A}$ anomaly effects lead to the so-called inverse mass hierarchy of $M_{\Xi_{Q}(1/2)^{-}}<M_{\Lambda_{Q}(1/2^{-})}$, or in terms of diquarks, $M_{[sq](0^{-})}<M_{[qq](0^{-})}$, despite their quark contents Harada et al. (2020); Suenaga and Oka (2023, 2024). QC2D also serves as a useful testing ground to explore possible diquark properties. The present LSM analysis can be straightforwardly extended to the $N_{f}=2+1$ case, where both [$qq$] and [$sq$] diquarks are included. Therefore, it is worthwhile to investigate the fate of the inverse mass hierarchy at finite temperature, particularly in relation to possible anomaly enhancements. This is left for future work.
@@ Acknowledgments
Appendix A Derivation of flow equations for $\lambda_{1k}$, $\lambda_{2k}$, $c_{1k}$ and $c_{2k}$
In Sec. III.3 it is demonstrated that the flow equations for $m_{k}$ and $a_{k}$ can be obtained via evaluating part of the effective potential that is quadratic in the field variables. In this Appendix, we show derivations of flow equations for the remaining quartic couplings: $\lambda_{1k}$, $\lambda_{2k}$, $c_{1k}$ and $c_{2k}$.
@@ Appendix A Derivation of flow equations for $\lambda_{1k}$, $\lambda_{2k}$, $c_{1k}$ and $c_{2k}$
where the definitions of ${\cal D}_{0,k}$ and $V_{(2)k}^{\prime\prime}$ are given by Eq. (38), see also the explanations below Eq. (LABEL:OmegaExpanded). As long as we stick to the zero chemical potential, ${\cal D}_{0,k}$ is diagonal and no Matsubara frequencies $\omega_{n}$ appear in any numerator, thus the evaluation of the RHS of Eq. (60) can be performed in a straightforward way. Making use of the identity
@@
$\partial_{k}\hat{V}_{B,k}^{I_{A}^{2}}$ $=$ $-\frac{1}{4}\int\frac{d^{3}p}{(2\pi)^{3}}\tilde{\partial}_{k}$
@@
$\partial_{k}\hat{V}_{B,k}^{I_{1}I_{A}}=-\frac{1}{4}\int\frac{d^{3}p}{(2\pi)^{3}}\tilde{\partial}_{k}\Bigg{(}\frac{(4c_{1,k}+2c_{2,k}-\lambda_{2,k})(2c_{2,k}+4\lambda_{1,k}+\lambda_{2,k})}{2(E_{k}^{+})^{2}}F_{1}(E_{k}^{+})$
@@
are representing contributions of mesonic and baryonic fluctuations to the respective coefficients. In addition, we note that $E_{k}^{\pm}=\sqrt{{\bm{p}}^{2}+R_{k}({\bm{p}})+(m^{\pm}_{{\rm eff},k})^{2}}$ with $(m^{\pm}_{{\rm eff},k})^{2}=m_{k}^{2}\pm a_{k}$, and ~$\tilde{\partial}_{k}$ is a differential operator acting only on the regulator $R_{k}$. In the above equations we have also defined
@@
Using Eqs. (64) - (70), flow equations for the quartic couplings are obtained by
@@
The ~$\tilde{\partial}_{k}$ derivatives are evaluated using
@@
×$\ \ \ \ \times\sum_{\alpha=\pm}\left[\frac{k}{E_{k}^{\alpha}}\frac{\partial{\cal F}_{k}(E_{k}^{+},E_{k}^{-})}{\partial E_{k}^{\alpha}}\right]_{E_{k}^{\pm}=\sqrt{k^{2}+(m^{\pm}_{{\rm eff},k})^{2}}}$

### v2

@@ IV.3 Enhancement of anomaly effects at finite $T$
For any values of $\mathring{M}_{\eta}$, all the coefficients exhibit significant increase at finite temperature; that is, the anomaly enhancement in hot medium is predicted to be similar in QC2D to the case of (three-color) QCD Fejos:2016hbp.
The enhancement of the quadratic anomaly coefficient $a$ is milder than the quartic ones, $c_{1}$ and $c_{2}$, where a larger $\mathring{M}_{\eta}$ leads to a smaller increase. For $c_{1}$ and $c_{2}$, the same ordering of the enhancement is seen up to $T\sim 0.2$ GeV, while at higher temperatures it is flipped. In general, the coefficients show a significant increase at low $T$, while at high $T$ the enhancements become tempered and begin to decrease. Note that the inversions occur when the chiral condensate starts to reduce (see Fig. 4). In addition, such nonmonotonic behaviors appear at lower $T$ for smaller values of $\mathring{M}_{\eta}$. We also note that $a<0$, $c_{1}<0$, and $c_{2}>0$ at any $T$.
@@ IV.4 Topological susceptibility
In terms of the fermionic language, this quantity can be reexpressed through two-point functions of pseudoscalar operators $\bar{\psi}i\gamma_{5}\psi$ and $\bar{\psi}i\gamma_{5}\tau_{f}^{a}\psi$, by the virtue of axial Ward-Takahashi identities GomezNicola:2016ssy; Kawaguchi:2020qvg. Then, within the LSM, imposing matching conditions appropriately, the topological susceptibility is evaluated as (for the detailed derivation, see Ref. Kawaguchi:2023olk)
@@
are pion and $\eta$-meson propagators, respectively, in momentum space.
In the present work, $D_{\pi}$ and $D_{\eta}$ are always simply expressed by $D_{\pi}=i/(p^{2}-M_{\pi}^{2})$ and $D_{\eta}=i/(p^{2}-M_{\eta}^{2})$, respectively. Note that their functional forms do not differ even at finite $T$, owing to the absence of diquark condensates Kawaguchi:2023olk. Hence, the susceptibility reads
@@
where we have defined the dimensionless topological susceptibility $\chi_{\rm top}$, normalized by the constant $\mathring{M}_{\pi}^{2}\mathring{\sigma}_{0}^{2}$. The mass difference between the $\eta$ meson and the pion is generated solely by anomaly effects [see Eq. (84)]; therefore, it is beneficial to define
@@
Depicted in Fig. 5 are the resulting $T$ dependencies of the normalized topological susceptibilities $\chi_{\rm top}$ and $\tilde{\chi}_{\rm top}$ with $\mathring{M}_{\eta}=0.3$, $0.5$, and $0.95$ GeV. Differences between the solid and dashed curves are always tiny, meaning that the anomaly enhancement effects do not generate sizable contributions to the topological susceptibility. Meanwhile, the dependence of the $\eta$ meson mass on the topological susceptibility is manifest only for larger $\mathring{M}_{\eta}$, which increases the magnitude of $\chi_{\rm top}$. Up to $T\sim 0.1$ GeV, the susceptibilities are almost constant for any values of $\mathring{M}_{\eta}$, while above this temperature they start to reduce. This reduction is understood by the smooth chiral symmetry restoration exhibited in Fig. 4, since the $T$ dependence of the susceptibility (58) is governed by
@@
Figure 6: $k$ dependencies of the anomaly coefficients $c_{1,k}$ and $c_{2,k}$ evaluated by the full (solid lines) and truncated (dashed lines) analyses at $T=0$, with $\mathring{M}_{\eta}=0.5$ GeV. Note that the UV scale is set to be $\Lambda_{\rm UV}=1$ GeV.
@@
Figure 8: $T$ dependencies of $\Delta\tilde{a}$, $\Delta\tilde{c}_{1}$, and $\Delta\tilde{c}_{2}$ defined in Eq. (61), with $\mathring{M}_{\eta}=0.5$ GeV.
@@ V Importance of baryonic contributions
In the previous section, modifications of the hadron masses and anomaly coefficients were examined. In this section, we shed light on the role of baryonic contributions in the FRG framework to gain insight into fluctuations specific to QC2D, where diquark baryons are also counted as hadrons. Here, we concentrate on the case where $\mathring{M}_{\eta}=0.5$ GeV.
In order to study baryonic contributions, on top of the full solution, we also solve the flow equations containing solely mesonic contributions (e.g., $\partial_{k}m^{2}_{k}=\tilde{\partial}_{k}\hat{V}^{I_{1}}_{M,k}$ for $m^{2}_{k}$) with inputs of $\mathring{M}_{\pi}=0.14$ GeV, $\mathring{M}_{\eta}=0.5$ GeV, $\lambda_{1,\Lambda}=\lambda_{2,\Lambda}/4$ and $c_{1,\Lambda}=c_{2,\Lambda}=0$. The value of $\lambda_{1,\Lambda}$ is adjusted such that $\mathring{M}_{\sigma}\approx 0.18$ GeV. Then, we compare the cases with and without baryons.
The parameters $m^{2}_{k}$, $a_{k}$ and $\lambda_{1,k}$ at $k=0$ are roughly determined by the masses of the pion, $\eta$, and $\sigma$ mesons, respectively, and, hence those values in the truncated and in the full analysis do not differ much. On the other hand, $c_{1,k}$ and $c_{k,2}$ are simply fixed to vanish at $k=\Lambda_{\rm UV}$, and, thus, their scale evolutions are expected to be a useful measure to see the differences between the two approaches. Depicted in Fig. 6 are the resultant $k$ dependencies for $c_{1,k}$ and $c_{2,k}$ at $T=0$, evaluated separately for the full solution and that containing only mesonic contributions. This figure clearly shows that the evolutions of $c_{1,k}$ and $c_{2,k}$ start to deviate in the two cases, soon after departing from $k=\Lambda_{\rm UV}$. As a result, at $k=0$, the coefficients containing both mesonic and baryonic contributions are significantly magnified compared to the truncated case, indicating the important role of baryonic fluctuations. Such enhancements can be simply understood by an increase in the degrees of freedom in the flows.
Finally, we also depict in Figs. 7 and 8 the temperature dependencies of the anomaly coefficients at $k=0$, evaluated using the full and the truncated analyses. In Fig. 8, for instance, $\Delta\tilde{a}$ is defined by
$\Delta\tilde{a} \displaystyle\equiv$ $\frac{\Delta a}{\Delta a|_{T=0}}\ ,$ (61)
where
$\Delta a=a|_{\rm meson+baryon}-a|_{\rm meson},$ (62)
measuring differences with and without baryonic contributions. The same applies to $\Delta\tilde{c}_{1}$ and $\Delta\tilde{c}_{2}$. These figures also show the importance of baryonic contributions at any temperature with regards to the anomaly coefficients.
@@ VI Conclusions
In this paper, we have investigated modifications of $U(1)_{A}$ anomaly effects at finite temperature in QC2D with two flavors, by means of the FRG technique. We have adopted the LSM developed in Ref. Suenaga:2022uqn to properly take into account fluctuations of the low-lying scalar and pseudoscalar mesons, as well as baryons, based on a linear representation of the Pauli-Gürsey $SU(4)$ symmetry.
As a result, we have found that quartic couplings of the LSM need to be suppressed, in order to properly reproduce a reasonable reduction of the chiral condensate at finite temperature. Accordingly, mass differences between the chiral partners $\sigma$ and $\pi$, or $a_{0}$ and $\eta$, are always found to be tiny, regardless of the $U(1)$ axial anomaly strength, i.e., the mass of the $\eta$ meson. At higher temperatures, mass degeneracies of the chiral partners in accordance with chiral symmetry restoration are realized. We note that masses of the $0^{+}$ and $0^{-}$ (anti)diquark baryons degenerate with those of $\pi$ and $a_{0}$, respectively, owing to the Pauli-Gürsey $SU(4)$ symmetry. The topological susceptibility at high temperature is found to be suppressed due to the reduction of chiral condensate.
Once we enter the finite-temperature regime, all of the couplings responsible for the $U(1)$ axial anomaly become magnified, and, hence we have concluded that due to hadronic fluctuations, the axial anomaly is enhanced in a hot medium for QC2D. Similar anomaly enhancements were also discussed in three-color QCD analyses Fejos:2016hbp. At high temperatures, the increase in the quartic couplings is gradually suppressed and eventually diminishes. Our results also confirm that baryonic fluctuations play a substantial role in the enhancements.
@@
We note that, while the $U(1)_{A}$ anomaly effects are enhanced at finite temperature within the hadronic model, LSM, it is known that at sufficiently high temperature and/or density, instantons melt due to the Debye screening of electric gluons, for which the $U(1)_{A}$ symmetry is effectively restored Gross:1980br; Rapp:1999qa. In order to smoothly connect those two behaviors, it would be necessary to implement a crossover transition to quark matter, but it is beyond the scope of the present study.
@@
Our analysis can also be applied to the case of a finite chemical potential. In the low-temperature and high-density regime, the emergence of a baryon superfluid phase, characterized by the condensation of positive-parity diquarks, has been demonstrated in lattice QCD simulations Boz:2019enj; Buividovich:2020dks; Astrakhantsev:2020tdl; Iida:2024irv; Braguta:2023yhd and model analyses Kogut:2000ek; Ratti:2004ra; Sun:2007fc; Suenaga:2022uqn; Suenaga:2025sln. In this phase, in particular, the possibility of enhanced $U(1)_{A}$ anomaly effects has been discussed by comparing the lattice data with LSM analyses of the hadron mass spectrum and topological susceptibilities  Suenaga:2022uqn; Kawaguchi:2023olk. Hence, investigating the fate of the $U(1)_{A}$ anomaly effects in the superfluid phase is considered an important direction for future work.
Finally, we provide some comments regarding diquarks. In three-color QCD, the chiral dynamics of diquarks is reflected in the spectrum of singly heavy baryons, which consist of a heavy quark $Q$ and a light diquark, either [$qq$] or [$sq$] ($q=u,d$). In this context, it is suggested that the $U(1)_{A}$ anomaly effects lead to the so-called inverse mass hierarchy of $M_{\Xi_{Q}(1/2)^{-}}<M_{\Lambda_{Q}(1/2^{-})}$, or in terms of diquarks, $M_{[sq](0^{-})}<M_{[qq](0^{-})}$, despite their quark contents Harada:2019udr; Suenaga:2023tcy; Suenaga:2024vwr. QC2D also serves as a useful testing ground to explore possible diquark properties. The present LSM analysis can be straightforwardly extended to the $N_{f}=2+1$ case, where both [$qq$] and [$sq$] diquarks are included. Therefore, it is worthwhile to investigate the fate of the inverse mass hierarchy at finite temperature, particularly in relation to possible anomaly enhancements. This is left for future work.
@@ Acknowledgments
Appendix A Derivation of flow equations for $\lambda_{1k}$, $\lambda_{2k}$, $c_{1k}$, and $c_{2k}$
In Sec. III.3, it is demonstrated that the flow equations for $m_{k}$ and $a_{k}$ can be obtained via evaluating part of the effective potential that is quadratic in the field variables. In this appendix, we show derivations of flow equations for the remaining quartic couplings: $\lambda_{1k}$, $\lambda_{2k}$, $c_{1k}$, and $c_{2k}$.
@@ Appendix A Derivation of flow equations for $\lambda_{1k}$, $\lambda_{2k}$, $c_{1k}$, and $c_{2k}$
where the definitions of ${\cal D}_{0,k}$ and $V_{(2)k}^{\prime\prime}$ are given by Eq. (39); see also the explanations below Eq. (LABEL:OmegaExpanded). As long as we stick to the zero chemical potential, ${\cal D}_{0,k}$ is diagonal and no Matsubara frequencies $\omega_{n}$ appear in any numerator; thus, the evaluation of the rhs of Eq. (63) can be performed in a straightforward way. Making use of the identity
@@
$\partial_{k}\hat{V}_{B,k}^{I_{A}^{2}}$ $=$ $-\frac{1}{4}\int\frac{d^{3}p}{(2\pi)^{3}}\tilde{\partial}_{k}$ (72)
@@
$\hskip-128.0374pt\partial_{k}\hat{V}_{B,k}^{I_{1}I_{A}}=-\frac{1}{4}\int\frac{d^{3}p}{(2\pi)^{3}}\tilde{\partial}_{k}\Bigg(\frac{(4c_{1,k}+2c_{2,k}-\lambda_{2,k})(2c_{2,k}+4\lambda_{1,k}+\lambda_{2,k})}{2(E_{k}^{+})^{2}}F_{1}(E_{k}^{+})$
@@
are representing contributions of mesonic and baryonic fluctuations to the respective coefficients. In addition, we note that $E_{k}^{\pm}=\sqrt{{\bm{p}}^{2}+R_{k}({\bm{p}})+(m^{\pm}_{{\rm eff},k})^{2}}$ with $(m^{\pm}_{{\rm eff},k})^{2}=m_{k}^{2}\pm a_{k}$, and $\tilde{\partial}_{k}$ is a differential operator acting only on the regulator $R_{k}$. In the above equations, we have also defined
@@
Using Eqs. (67)–(74), flow equations for the quartic couplings are obtained by
@@
The $\tilde{\partial}_{k}$ derivatives are evaluated using
@@
$\ \ \ \ \times\sum_{\alpha=\pm}\left[\frac{k}{E_{k}^{\alpha}}\frac{\partial{\cal F}_{k}(E_{k}^{+},E_{k}^{-})}{\partial E_{k}^{\alpha}}\right]_{E_{k}^{\pm}=\sqrt{k^{2}+(m^{\pm}_{{\rm eff},k})^{2}}}$